# ARS: Automatic Routing Solver with Large Language Models

### v1

@@ Appendix C Prompts of ARS
Automatic Routing Solver (ARS) is designed to perform three key steps leveraging the capabilities of LLMs: Constraint Selection, Constraint Checking Program Generation, and Constraint Scoring Program Generation. In this section, we describe the prompt engineering involved in each step. These prompts are constructed based on user inputs and several representative constraints stored in the database to generate the Constraint Checking Program and the Constraint Scoring Program. Variable information, such as user inputs and constraint descriptions, is highlighted in blue for clarity.
@@
Step 1: Constraint Selection. In the first step, the ARS identifies constraints relevant to the user’s input from the database. This step processes the user input and matches it against the constraints stored in the database. If relevant constraints are found, they are selected for further processing. Otherwise, LLM outputs “No Relevant Constraint”. This step ensures that only the constraints pertinent to the user’s query are considered in subsequent steps.
@@
Prompt for Constraint Selection For the description in the VRP problem, identify and provide the relevant constraint types from the following list: {constraint_description} According to the user input: {user_input} If no constraint types match the user input, respond with: ”No Relevant Constraint”. Do not give additional explanations.
@@
Step 2: Constraint Checking Program Generation. Based on the relevant constraints selected in the previous step, the ARS uses the LLM to generate new constraints by taking the selected constraints as references. Specifically, the selected constraints are provided as input to the LLM, which then generates new constraints tailored to the user’s input.
@@
Step 3: Constraint Scoring Program Generation. In the final step, the ARS generates a Constraint Scoring Program based on the Constraint Checking Program created in the previous step. This scoring program evaluates the degree to which the constraints are satisfied, providing a quantitative measure of constraint adherence.
@@
Appendix D VRP Variants
@@
The common VRP variants may involve several representative constraints, such as vehicle capacity, distance limits, time windows, pickup and delivery, same vehicle, and priority. As shown in Table 6, these six representative constraints can be combined to generate 48 common problems, with the exception of cases where vehicle capacity and pickup and delivery appear together.
@@
Appendix E Experimental Details
Experiments are performed on a computer with an Intel Xeon Gold 6248R Processor (3.00 GHz), 128 GB system memory, and Windows 10.
@@ Appendix E Experimental Details
Appendix F Examples of Program Codes
@@
To better analyze different methods for solving VRPs, we provide code examples for three approaches: our solver, Gurobi, and OR-Tools. As a case study, these methods are applied to the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) to illustrate their respective requirements and complexities.
From the provided code examples, it is clear that our solver requires the least amount of code to be generated by an LLM. This is attributed to the simplicity and flexibility of our method, which avoids the need for verbose or overly rigid programming constructs. In contrast, the code examples for Gurobi and OR-Tools are more extensive, primarily due to their reliance on strict syntax rules and detailed configurations.
@@ 3 Args:
6 - "edge_weight": A 2D NumPy array of distances between nodes.
@@ 16 Note:
6 Finds feasible routes for a CVRP problem using an optimization solver.
@@ 16 Returns:
28 model = Model("CVRP_PathLength")
@@
65 (arrival_time[i] + service_time[i] + edge_weight[i][j] <= arrival_time[j] + (1 - x[i, j]) * 1e6 for i in
66 range(num_nodes) for j in range(1, num_nodes) if i != j), name="TravelTime")
@@ 5 Args:
9 - "demand": A 1D NumPy array of node demands (0 for the depot).
10 - "time_window": A 2D NumPy array where each row represents the [earliest start, latest end] time windows for servicing each node. The maximum time limit for each route is set to the depot’s latest end time.(0 for the depot)
13 Returns:
@@ 16 Note:
56 demand_callback_index,
57 0, # null capacity slack
58 [vehicle_capacity] * num_vehicles, # vehicle maximum capacities
59 True, # start cumul to zero
60 ’Capacity’
@@ 59 True, # start cumul to zero
73 time_callback_index,
74 horizon, # allow waiting time
75 horizon, # maximum time per vehicle
76 False, # don’t force start cumul to zero
77 ’Time’

### v3

@@ B.1 Prompts of ARS
Automatic Routing Solver (ARS) is designed to address each VRP variant by leveraging LLMs in three key steps: Constraint Selection, Constraint Checking Program Generation, and Constraint Scoring Program Generation, with a total of three LLM calls. In this section, we describe the prompt engineering involved in each step. These prompts are constructed based on user inputs and several representative constraints stored in the database to generate the Constraint Checking Program and the Constraint Scoring Program. Variable information, such as user inputs and constraint descriptions, is highlighted in blue for clarity.
@@
Step 1: Constraint Selection. In the first step, the ARS identifies constraints relevant to the user’s input from the database. This step processes the user input and matches it against the constraints stored in the database. If relevant constraints are found, they are selected for further processing. Otherwise, LLM outputs "No Relevant Constraint". This step ensures that only the constraints relevant to the problem description are considered in subsequent steps.
@@
Prompt for Constraint Selection For the description in the VRP problem, identify and provide the relevant constraint types from the following list: {constraint_description} According to the user input: {user_input} If no constraint types match the user input, respond with: "No Relevant Constraint". Do not give additional explanations.
@@
Step 2: Constraint Checking Program Generation. Based on the relevant constraints selected in the previous step, the ARS uses the LLM to generate a new Constraint Checking Program by taking the selected constraints as references. Specifically, the selected constraints are provided as input to the LLM, which then generates the new program tailored to the problem description and the referenced constraint information.
@@
Step 3: Constraint Scoring Program Generation. In the final step, the ARS generates a Constraint Scoring Program using the Constraint Checking Program developed in the previous step as a foundation. This scoring program evaluates the degree to which the constraints are satisfied by assigning a quantitative score based on the results of the constraint checks.
@@
B.2 Examples of LLM Outputs
This subsection shows an example of using LLM with ARS to solve the CVRP with Incompatible Loading Constraints (CVRP-ILS). The process includes three LLM calls. The first call selects constraints related to CVRP-ILS. The second call creates a program to check these constraints. The third call creates a program to calculate violation scores.
Example of Step 1: Constraint Selection First Call Input:
For the description in the VRP problem, identify and provide the relevant constraint types from the following list:
1. No Relevant Constraint Example: No relevant constraint. 2. Vehicle Capacity Constraint Example: The total load of each route must not exceed the vehicle capacity. 3. Distance Limit Constraint Example: Each route’s total length must not exceed 100 units. 4. Time Windows Constraint Example: The arrival time at the node must meet the time windows. 5. Pickup and Delivery Constraint Example: At Node [24], 10 units of goods are picked up. Unlike delivery, visiting a pickup node reduces the vehicle’s available capacity. 6. Same Vehicle Constraint Example: Nodes [1, 10] must be served by the same vehicle. 7. Priority Constraint Example: Nodes [17, 19] are priority points. According to the user input: The total load of each route must not exceed the vehicle capacity. Nodes [7, 8] must not be on the same route. If no constraint types match the user input, respond with: "No Relevant Constraint". Do not give additional explanations. First Call Output:
————————– Relevant Constraints Type ————————–
1. Constraint type: Vehicle Capacity Constraint 2. Constraint type: Same Vehicle Constraint
The problem description for CVRP-ILS is: The total load of each route must not exceed the vehicle capacity. Nodes [7, 8] must not be on the same route.
[…]
Example of Step 2: Constraint Checking Program Generation Second Call Input:
As a Python algorithm expert, please implement a function to check the constraints for the vehicle routing problem (VRP) based on the provided description and relevant code.
User input: The total load of each route must not exceed the vehicle capacity. Nodes [7, 8] must not be on the same route. Relevant Examples: # See Appendices G.2 and G.6. Do not give additional explanations. Second Call Output: def check_constraints(solution: VrpState) -> bool: """ Check if the solution satisfies the constraints. Args: solution (VrpState): An object representing the VRP solution. - problem_data (dict): A dictionary with problem details: - "edge_weight": A 2D NumPy array of node distances . - "demand": A 1D NumPy array of node demands (0 for the depot). - "capacity": Maximum load capacity per vehicle. Return: bool: True if the solution satisfies the constraints, False otherwise. Note: The above "Args" cannot be added or modified, and no data outside the "solution" object should be used or added. """ # Check vehicle capacity constraint for route in solution.routes: total_demand = sum(solution.problem_data["demand"][node] for node in route) if total_demand > solution.problem_data["capacity"]: return False # Check nodes [7, 8] not on same route constraint for route in solution.routes: if 7 in route and 8 in route: return False return True
Example of Step 3: Constraint Scoring Program Generation Third Call Input:
As a Python algorithm expert, please implement a function to calculate the constraint violation score for the Vehicle Routing Problem (VRP) based on the given constraints.
Function Template: # See Appendix F. Constraints Description: The total load of each route must not exceed the vehicle capacity. Nodes [7, 8] must not be on the same route. Check Constraints Code: # Use the constraint checking code generated in Step 2. Do not give additional explanations. Third Call Output: def calculate_violation_score(solution: VrpState) -> float: """ Args: solution (VrpState): An object representing the VRP solution. - problem_data (dict): A dictionary with problem details; Use only defined values in "Check Constraints Code". - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot). - cost (Callable[[], float]): Implemented method to compute and return the total cost of the solutions routes. Return: violation_score (float): The calculated constraint violation score, where higher scores indicate more severe violations. Note: Do not use data or methods that is not contained in the "solution" object. """ violation_score = 0.0 # Check vehicle capacity constraint for route in solution.routes: total_demand = sum(solution.problem_data["demand"][node] for node in route) if total_demand > solution.problem_data["capacity"]: violation_score += (total_demand - solution.problem_data["capacity"]) # Check nodes [7, 8] not on same route constraint for route in solution.routes: if 7 in route and 8 in route: violation_score += 1.0 return violation_score
B.3 Operator
Local search operators are essential components of heuristic and metaheuristic methods, designed to explore the neighborhood of a solution and iteratively improve its quality. These operators are the building blocks for efficiently navigating the search space, balancing exploration and exploitation. The commonly used operators are as follows:
2-opt Operator [24]. The 2-opt operator is a classical approach originally developed for the Traveling Salesman Problem (TSP). It works by removing two non-adjacent edges in the solution and reconnecting them in a different way, thereby altering the order of nodes. If the new configuration reduces the total cost, it is accepted as an improved solution.
Swap Operator [25]. The Swap operator is another simple yet powerful tool in local search methods. It works by exchanging the positions of two elements within the solution. This operation generates a new configuration, which can help in escaping local optima and promoting diversity in the solution space.
Shift Operator [26]. The Shift operator involves moving an element from one position in the solution to another. This operation changes the relative ordering of elements, redistributing their positions to explore alternative configurations. By shifting elements, the algorithm can adjust the structure of the solution in a more targeted manner, allowing it to overcome local optimality and discover new regions of the solution space.
Destroy Operator. The Destroy operator partially disrupts the current solution by selectively removing some elements. This disruption breaks the local optimality of the solution, allowing for the exploration of new regions in the search space. There are two common implementations of this operator: Random Removal and String Removal.
  • Random Removal: This method involves removing elements uniformly at random, without relying on specific heuristics or problem-dependent strategies, making it a straightforward yet highly effective approach to diversify the search process and introduce variability into the solution space.
  • String Removal: This method targets sequences of consecutive or related elements (strings), such as partial routes or groups of customers [23]. It begins by randomly selecting a "center" customer and removing a string of nearby customers from the route. The string size is randomly determined, constrained by the average route size and a predefined maximum. If constraints on the number of disrupted routes or previously disrupted routes are met, further removal is skipped.
Repair Operator. The Repair operator complements the Destroy operator by reinserting removed elements to reconstruct a complete solution, guided by optimization objectives. This combination of destruction and repair allows the algorithm to iteratively refine solutions while maintaining the flexibility to explore new possibilities. One commonly used implementation of the Repair operator is Greedy Repair:
  • Greedy Repair: This method reinserts removed elements one by one, selecting at each step the position that minimizes the objective function. By considering constraint-aware heuristics during the reinsertion process, it ensures that each step improves solution quality while adhering to problem-specific constraints, effectively balancing optimality and feasibility throughout the search process.
In summary, the local search operators discussed above, including 2-opt, Swap, Shift, Destroy, and Repair, play a crucial role in the design of heuristic and metaheuristic algorithms. These operators enable targeted adjustments to the solution, facilitating efficient exploration and exploitation of the solution space. By combining these operators, algorithms can effectively escape local optima and converge toward high-quality solutions.
In our approach, we adopt these efficient operators within a Backbone heuristic framework, which provides the structural foundation for solving complex optimization problems. The framework leverages these operators to iteratively refine solutions, balancing between intensification and diversification.
Appendix C VRP Variants
@@
Common problems in VRP are typically constructed based on a set of fundamental constraints, such as vehicle capacity, distance limits, time windows, pickup and delivery, same vehicle, and priority. These problems are widely used to evaluate the performance of multi-task algorithms [20, 73, 75]. By combining six representative constraints, excluding cases where vehicle capacity conflicts with pickup and delivery, a subset of 48 common problems is generated, as shown in Table 5.
@@
Appendix D Experimental Details
Experimental environment: Experiments are performed on a computer with an Intel Xeon Gold 6248R Processor (3.00 GHz), 128 GB system memory, and Windows 10.
D.1 Analysis of ARS in RoutBench
@@
D.2 The Number of LLM-Suggested Constraints
To better understand the relationship between the real constraint count in VRPs and the constraints suggested by the LLM agent, we analyze how the LLM adapts its recommendations based on the problem’s complexity. As shown in Table 7, the data reveals a clear trend where the number of LLM-suggested constraints increases as the real constraint count grows. This suggests that the LLM agent effectively adapts its recommendations based on the complexity of the problem. Interestingly, the LLM agent tends to suggest slightly more constraints than the actual count, likely as a precautionary measure to ensure no potentially relevant constraints are overlooked.
Table 7: Analysis of the relationship between real constraint counts and LLM-suggested constraints in various VRPs, including the total number of problems analyzed for each constraint count and the corresponding average number of suggested constraints by the LLM agent.
Constraint Count Number of Problems Average of LLM-Suggested Constraints
1 8 1.25
2 92 2.43
3 400 3.81
4 263 4.84
5 237 5.63
D.3 Enhancing ARS with Other Methods
To further improve the performance of ARS in program generation tasks, we enhanced it with two methods: Reflexion [61] and Self-debug [29]. Table 8 presents a detailed comparison of the original ARS and its enhanced variants (ARS+Reflexion and ARS+Self-debug) on RoutBench. The results, obtained using the DeepSeek V3, show that both enhancements lead to improvements in success rate (SR) while simultaneously reducing the error rate (RER). Specifically, ARS+Self-debug achieves the highest success rates and the lowest error rates on RoutBench, demonstrating its effectiveness in refining the program generation process. These findings highlight the potential of incorporating self-improvement mechanisms to boost the performance of ARS.
Table 8: Performance comparison of ARS and its enhanced variants (ARS+Reflexion and ARS+Self-debug) on RoutBench. The best results among these methods are highlighted in grey.
Methods RoutBench-S RoutBench-H
SR $\uparrow$ RER $\downarrow$ SR $\uparrow$ RER $\downarrow$
ARS+Reflexion 78.20% 0.20% 63.20% 2.00%
ARS+Self-debug 78.80% 0.00% 68.80% 0.60%
ARS 77.20% 2.80% 61.60% 5.60%
D.4 Performance of ARS across Multiple Runs
We conducted three independent runs of ARS to evaluate its performance on program generation tasks, with the detailed results presented in Table 9. The experiments are performed on RoutBench using the DeepSeek V3, reporting success rates (SR) and error rate reductions (RER) for each trial, along with the averages and standard deviations. The results show consistent performance across the three runs, with low standard deviations in both SR and RER, indicating that ARS performs stably and reliably across different runs and benchmarks.
Table 9: Detailed results of three independent runs of ARS on RoutBench, including success rates, runtime error rates, averages, and standard deviations to evaluate stability and reliability.
Experiment RoutBench-S RoutBench-H
SR $\uparrow$ RER $\downarrow$ SR $\uparrow$ RER $\downarrow$
Trial 1 77.20% 2.80% 61.60% 5.60%
Trial 2 80.60% 2.00% 64.20% 2.60%
Trial 3 76.40% 2.80% 64.80% 5.80%
Average 78.07% 2.53% 63.53% 4.67%
Standard Deviation 3.56% 0.14% 2.14% 2.89%
D.5 Best-Known Solutions for RoutBench
The best-known solutions (BKS) can be accessed through the provided link to the RoutBench repository. We provide the BKS for all instances in RoutBench, a comprehensive benchmark that encompasses 1,000 VRP variants with varying problem sizes (25, 50, and 100 nodes). For each instance, the BKS is obtained using ARS, which applies the correct Constraint-Aware Heuristic to ensure feasibility and solution quality. The algorithm is executed under rigorous stopping criteria: it terminates when no improvement is observed for 1,000 consecutive generations or when the runtime exceeds one hour.
D.6 Problem Set for Test Instances
As shown in Table 10, two common problems (e.g., CVRP and CVRPTW) and two dynamic problems (e.g., DCVRP and DCVRP-L) are presented, along with their respective settings and constraints. In this paper, "N" represents the number of points, "C" represents the vehicle capacity, and "L" represents the length of the maximum travel distance.
Table 10: Detailed problem set and constraints for test instances, including common and dynamic VRP variants.
Problem Problem Set Problem Description
CVRP C=200 The total load of each route must not exceed the vehicle capacity.
CVRPTW C=200 The total load of each route must not exceed the vehicle capacity. The arrival time at each node must meet its specified time window.
DCVRP C=200 The total load of each route must not exceed the vehicle capacity. Specifically, for node [19], its base demand is augmented by 5 times the square root of the accumulated travel distance from the depot [0] to that node.
DCVRP-L C=200, L=150 The total load of each route must not exceed the vehicle capacity. Specifically, for node [19], its base demand is augmented by 5 times the square root of the accumulated travel distance from the depot [0] to that node. Each route must not exceed 150 units in length.
D.7 Evaluation of ARS Performance on CVRPLib
To further validate the effectiveness of ARS in solving real-world instances, we conducted experiments using five test suites from the CVRPLib [^2] http://vrp.atd-lab.inf.puc-rio.br/ benchmark datasets. These datasets consist of 99 instances from Sets A, B, F, P, and X [80], encompassing graph scales ranging from 16 to 200 nodes, diverse node distributions, and varying customer demands.
In this subsection, we provided benchmark results for state-of-the-art algorithms HGS and LKH-3, as well as some simple heuristics, including the Greedy Algorithm and Hill Climbing, for reference purposes. Additionally, we compared ARS with three widely-used solvers: CPLEX, OR-Tools, and Gurobi.
As shown in Table 11, ARS achieves the best performance among the four solvers, demonstrating its effectiveness and reliability in solving these challenging instances. The data for HGS, LKH-3, and simple heuristics serve as a reference to contextualize the problem complexity and solution quality.
Table 11: Performance comparison of ARS with other methods on CVRPLib. The best results among the four solvers (e.g., CPLEX, OR-Tools, Gurobi, and ARS) are highlighted in grey.
HGS LKH-3 Greedy Hill Climbing
Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap
Set A 1041.9 1042.2 0.00% 1041.9 0.00% 1379.6 32.41% 1183 13.54%
Set B 963.7 964.5 0.00% 963.7 0.00% 1272 31.99% 1140.6 18.36%
Set F 707.7 709 0.00% 707.7 0.00% 1021 44.27% 1044.3 47.56%
Set P 587.4 586.9 0.00% 587.4 0.00% 728.7 24.06% 681 15.93%
Set X 27220.1 27223.7 0.01% 27281.4 0.02% 34254.3 25.84% 32301.3 18.67%
Avg. 6104.2 6105.3 0.00% 6116.4 0.00% 7731.1 31.71% 7270.1 22.81%
CPLEX OR-Tools Gurobi ARS
Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap
Set A 1041.9 1096.5 5.24% 1058.9 1.63% 1067.3 2.44% 1055.5 1.31%
Set B 963.7 1003.6 4.14% 973.3 1.00% 990.9 2.82% 973 0.96%
Set F 707.7 789.3 11.53% 728.7 2.97% 728.7 2.97% 727 2.73%
Set P 587.4 612.5 4.27% 592 0.78% 594.9 1.28% 591.1 0.62%
Set X 27220.1 32044.1 17.72% 28209.6 3.64% 28977.7 6.46% 28142.4 3.39%
Avg. 6104.2 7109.2 8.58% 6312.5 2.00% 6471.9 3.19% 6297.8 1.80%
Appendix E Examples of Solver Codes
@@
To better analyze different methods for solving VRPs, we provide code examples for four approaches: our solver, Gurobi, OR-Tools, and CPLEX. As a case study, these methods are applied to the Capacitated Vehicle Routing Problem with Time Windows (CVRPTW) to illustrate their respective requirements and complexities.
@@ 3 Args:
6 - "edge_weight": A 2D NumPy array of node distances.
@@ 16 Note:
6 Finds feasible routes for a CVRPTW using an optimization solver.
@@ 16 Return:
28 model = Model("CVRPTW_PathLength")
@@
66 (arrival_time[i] + service_time[i] + edge_weight[i][j] <= arrival_time[j] + (1 - x[i, j]) * 1e6
67 for i in range(num_nodes) for j in range(1, num_nodes) if i != j), name="TravelTime")
@@
From the provided code examples, it is clear that our solver requires the least amount of code to be generated by an LLM. This is attributed to the simplicity and flexibility of our method, which avoids the need for verbose or overly rigid programming constructs. In contrast, the code examples for Gurobi, OR-Tools, and CPLEX are more complex, primarily due to their reliance on strict syntax rules and detailed configurations.
@@ 5 Args:
9 - "demand": 1D numpy array, demand [0]=0.
10 - "time_window": A 2D NumPy array where each row represents the [earliest, latest] time windows for servicing each node.
11 - "service_time": service time for each node.
12 - "edge_weight": 2D distance matrix.
13 Return:
@@ 16 Note:
56 demand_callback_index, 0, [vehicle_capacity] * num_vehicles, True, ’Capacity’
@@
69 time_callback_index, horizon, horizon, False, ’Time’
@@
CPLEX Code:
1from docplex.mp.model import Model
2
3def find_feasible_routes(solution: VrpState) -> list[list[int]]:
4 """
5 Finds feasible routes for a CVRPTW using CPLEX/DOcplex.
6
7 Args:
8 solution (VrpState): Contains problem_data dict with keys:
9 - "capacity": Q
10 - "demand": 1D numpy array, demand[0]=0
11 - "edge_weight": 2D distance matrix
12 - "time_window": time window, time_window[i]=(earliest, latest)
13 - "service_time": service time for each node
14 - optionally "num_vehicles": K
15 Returns:
16 List of routes (each a list of customer indices, excluding depot 0).
17 """
18 data = solution.problem_data
19 Q = data["capacity"]
20 demand = data["demand"]
21 edge_weight = data["edge_weight"]
22 time_window = data["time_window"]
23 service_time = data["service_time"]
24 num_nodes = len(demand)
25 V = list(range(num_nodes))
26 N = V[1:]
27 K = data.get("num_vehicles", None)
28
29 # Get the earliest and latest times for time windows
30 earliest = {{i: time_window[i][0] for i in V}}
31 latest = {{i: time_window[i][1] for i in V}}
32
33 # Calculate a large number M as the upper bound for time constraints
34 M = sum(latest[i] + service_time[i] + max(edge_weight[i]) for i in V)
35
36 # Build model
37 model = Model(name="VRPTW")
38 model.parameters.timelimit = {run_time} # Set time limit for solving
39
40 # Decision variables x[i,j]
41 arcs = [(i, j) for i in V for j in V if i != j]
42 x = model.binary_var_dict(arcs, name="x")
43
44 # MTZ load variables u[i] with lb=demand[i], ub=Q
45 u = model.continuous_var_dict(
46 N,
47 lb={{i: float(demand[i]) for i in N}},
48 ub=Q,
49 name="u"
50 )
51 # Time variables: arrival time at node i
52 t = model.continuous_var_dict(V, lb=0, name="t")
53
54 # Objective: minimize total distance
55 model.minimize(model.sum(edge_weight[i, j] * x[i, j] for i, j in arcs))
56
57 # Degree constraints: each customer exactly one in and one out
58 for i in N:
59 model.add_constraint(model.sum(x[i, j] for j in V if j != i) == 1)
60 model.add_constraint(model.sum(x[j, i] for j in V if j != i) == 1)
61
62 # Vehicle count limit at depot
63 if K is not None:
64 model.add_constraint(model.sum(x[0, j] for j in N) <= K)
65 model.add_constraint(model.sum(x[i, 0] for i in N) <= K)
66
67 # MTZ sub-tour elimination / capacity constraints
68 for i, j in arcs:
69 if i != 0 and j != 0:
70 model.add_constraint(u[i] + demand[j] <= u[j] + Q * (1 - x[i, j]))
71
72 # Time window constraints
73 # Enforce time window ranges
74 for i in V:
75 model.add_constraint(earliest[i] <= t[i])
76 model.add_constraint(t[i] <= latest[i])
77
78 # Ensure time logic along the route is correct
79 for i, j in arcs:
80 if j != 0: # No need to consider time window for returning to the depot
81 model.add_constraint(
82 t[i] + service_time[i] + edge_weight[i, j] <= t[j] + M * (1 - x[i, j])
83 )
84
85 # Solve
86 sol = model.solve(log_output=True)
87 if sol is None:
88 logger.error("No solution found.")
89 return []
90
91 # Extract used arcs
92 used = [(i, j) for (i, j) in arcs if x[i, j].solution_value > 0.5]
93
94 # Build successor map
95 succ = {{i: j for i, j in used}}
96
97 # Reconstruct routes from depot
98 routes = []
99 starts = [j for i, j in used if i == 0]
100 for st in starts:
101 route = []
102 cur = st
103 while cur != 0:
104 route.append(cur)
105 cur = succ.get(cur, 0)
106 routes.append(route)
107
108 return routes
Appendix F Function Template
The Constraint Scoring Program is designed to guide the generation of scoring functions in a structured manner. The following code template serves as a framework to define the scoring function, ensuring clarity and proper parameter usage. Within the template, the function name is predefined, and the roles of relevant parameters are described in detail. The specific implementation details are generated by the LLM based on the target VRP variant.
1def calculate_violation_score(solution: VrpState) -> float:
2 """
3 Args:
4 solution (VrpState): An object representing the VRP solution.
5 - problem_data (dict): A dictionary with problem details; Use only defined values in "Check Constraints Code".
6 - routes (list): A list of routes, where each route is a list of node IDs (integers, excluding the depot).
7 - cost (Callable[[], float]): Implemented method to compute and return the total cost of the solution’s routes.
8 Returns:
9 violation_score (float): The calculated constraint violation score, where higher scores indicate more severe violations.
10
11 Note:
12 Do not use data or methods that is not contained in the "solution" object.
13 """
14 # Your code goes here ...